X and Y intercept for y=-x2-4x-3

Answer by mananth(16949)

y=-x2-4x-3

for y intercept plug x=0
y=-3
for x co ordinate plug y =0
-x^2-4x-3=0
-x^2-3x-x-3=0
-x(x+3)-1(x+3)=0
(x+3)(-x-1)=0
x=-3 OR x=-1